### v1

@@ 4.3 Autonomous Learning via Reinforcement Optimization
Supervised fine-tuning transfers edge-level judgment and question quality from a strong evaluator, but it only imitates the teacher on logged cases. Imitation is static: it is limited by dataset coverage, does not learn when not to ask, and cannot balance accuracy against constraints under a new scenario. The ask gate and addressee are discrete choices whose value is revealed only after downstream effects, so credit must flow across the chain. We therefore continue with reinforcement learning to adapt the factored controller in Eq. (2) to the constrained objective in Eq. (3), using shaped rewards that turn these trade-offs into learnable signals.
@@
Second, we enforce parsimony to keep clarifications scarce and budget-aware via a sliding counter $c_{t}$:
@@
Third, we stabilize outputs with a light format bonus that preserves schema and shortness:
@@
These signals are aggregated into an edge reward that embodies our “fix-early, fix-locally” principle,
@@
Together, Eqs. (10)–(11) align the controller with our objective: intervene only when a concise question can halt DG/RD/SC/CG propagation, avoid gratuitous turns, keep outputs well-formed, and ultimately solve the task.
Optimization: E-GRPO
Before the episode reaches its terminal agent, the final correctness reward $R$ is unavailable; By following the Group Relative Policy Optimization (GRPO) (DeepSeek-AI et al., 2025), in this prefix regime, we train exactly on the edge-level signal by applying a within-edge relative update that selects the best clarification using $r_{t}^{\mathrm{edge}}$. Once the episode terminates and $R$ is observed, we augment the same update with a global credit that allocates outcome utility back to each edge occurrence, thereby coupling local containment with end-task success.
For each edge $t$, we form a local advantage $A_{t}^{\mathrm{loc}}$ by the controller using $r_{t}^{\mathrm{edge}}$. The system-level signal is distributed to every token at time $t$ with nonnegative weights $w_{t}$ and an action-independent baseline $b$,
$A_{t}^{\mathrm{glob}}=w_{t}\,(R-b).$ (12)
@@
When $R$ is not yet available, the second term in Eq. (13) is simply absent, and training proceeds purely on edge-level selection; when $R$ arrives, it is injected uniformly through $A_{t}^{\mathrm{glob}}$, ensuring that minimal, successful clarifications remain aligned with end-to-end correctness while updates stay conservative via clipping and reference KL (where “KL” is the Kullback–Leibler divergence between the current edge policy and the SFT-initialized reference, acting as a trust-region regularizer that preserves the clarifier’s brevity and schema prior).
@@ 5 Experiments
We assess whether a lightweight, edge-level clarifier improves reliability under unchanged orchestration, how it trades accuracy against latency and cost, and why these effects arise. Our study centers on three research questions (RQ):
  • Effectiveness under fixed orchestration (RQ1): Does AgentAsk improve end-task performance across frameworks and datasets without altering existing conditions? (Results on Table 1.)
  • Efficiency at the accuracy frontier (RQ2): Where does AgentAsk sit on the efficiency frontier relative to a strong evaluator and baselines? (Results: Table 2, breakdowns in Appendix C.)
  • Mechanism and Robustness at the edge (RQ3): Which error types dominate and how often are they one-shot resolvable, and how stable are outcomes under edge-level controls? (Summaries: Figure 4; cases in Appendix B)
@@ 5.1 Experimental Setup
We evaluate AgentAsk on five public benchmarks spanning math reasoning (GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021b)), general-knowledge QA (MMLU (Hendrycks et al., 2021a)), and code generation (HumanEval (Chen et al., 2024a), MBPP (Austin et al., 2021)).Baselines include single-model prompting (IO, CoT (Wei et al., 2022b)) and four representative multi-agent frameworks (GPTSwarm (Zhuge et al., 2024), AFlow (Zhang et al., 2025b), MaAS Zhang et al. (2025a), MasRouter (Yue et al., 2025)).
For each framework, we consider three settings that isolate the effect of the clarifier while holding orchestration constant: (i) origin uses the public implementation; (ii) +GPT-5 substitutes the clarifier with GPT-5 (OpenAI, 2025); (iii) +AgentAsk plugs in our lightweight module. The LLM executor is GPT-4o-mini-0718 (OpenAI, 2024); +AgentAsk uses Qwen-3-4B (Yang et al., 2025) as the backbone. Accuracy for GSM8K, MATH, and MMLU is reported; Pass@1 for HumanEval and MBPP. Prompts are provided in the Appendix D.
Figure 4: (Left) Pareto frontier (x=latency, y=extra cost; bubble size=Acc), showing +AgentAsk nearing +GPT-5 at much lower overhead. (Middle) Error-Type distributions (DG/SC/RD/CG) across datasets. (Right) Sensitivity to window $H$ and penalty $\lambda_{\mathrm{sw}}$, highlighting a stable region near the default and the accuracy–efficiency trade-off.
@@ 5.2 RQ1: Effectiveness under fixed orchestration
[…]
Results and analysis. Table 1 reports accuracy/ Pass@1 on five benchmarks across four frameworks under origin, +GPT-5, and +AgentAsk. Relative to origin, +AgentAsk yields positive deltas in roughly $17/20$ cells with mean gains around +0.5+\!0.5 pp while orchestration remains unchanged. Gains appear in math (GSM8K, MATH), general knowledge (MMLU), and code (HumanEval, MBPP), indicating broadly useful, link-local clarifications. A small number of non-monotone cells coincide with datasets whose error mixture is less amenable to one-shot clarification (see RQ3 in Section 5.4). We can see that GPT-5 generally achieves the greatest improvement, but our AgentAsk approach also achieves considerable performance gains. However, their performance on other metrics, such as cost and latency, remains to be compared, and we will conduct subsequent experiments and analysis.
  Settings GSM8K
Accuracy Latency Extra Cost
@@
Table 2: Ablations. Experiment on different type of clarifier with the MasRouter@GSM8K.
5.3 RQ2: Efficiency at the accuracy frontier
Observation. AgentAsk approaches the heavy-evaluator ceiling at a fraction of the latency and cost, while surpassing origin with small overhead.
Results and analysis. On MasRouter@GSM8K (Table 2), substituting a strong evaluator maximizes accuracy (95.34%) but raises latency and cost to 134% and 38.0%. In contrast, AgentAsk (Qwen-3-4B, E–GRPO) attains 94.72% at 103% latency and 4.2% extra cost—31 pp and 33.8 pp lower than +GPT-5, respectively, with only 0.62 pp accuracy gap. The Llama-3.2-3B variant mirrors this trend. Thus, AgentAsk situates on a favorable Pareto frontier: good-enough accuracy under tight budgets, offering a pragmatic alternative when heavy evaluators are cost-prohibitive.
In the Pareto view (Figure 4 Left), +AgentAsk bubbles cluster near the accuracy envelope but remain close to the origin on latency/cost axes, indicating that most of the attainable accuracy can be realized without incurring heavy-evaluator spend.
Experimental results on other datasets and baselines have been presented in the appendix C (See Table 7, Table 8, and Table 9) and briefly analyzed.
@@ 5.4 RQ3: Mechanism and Robustness at edge
Observation. DG/SC dominate and are frequently one-shot fixable; RD/CG are rarer and harder to neutralize in a single turn, matching where gains are largest.
Results and analysis. Figure 4(Middle) summarizes clarification types. Pooled over events, Data Gap and Signal Corruption constitute the majority (e.g., 36.8%, 30.9%) and exhibit higher one-shot resolution (about 72.5%, 69.1%), which corresponds to the broad positive deltas in Table 1. Referential Drift and Capability Gap are less frequent (20.7%, 11.6%) and less often one-shot resolved (56.4%, 44.8%), aligning with smaller or occasional negative deltas on RD/CG-richer subsets, notably parts of MATH/MMLU. Qualitative case studies pair each raw trace with the triggered question (when/what/whom) and the corrected downstream step; these examples are shown in Appendix B, making the edge-containment mechanism concrete.
Sensitivity. We further examine how two key parameters shape the accuracy–efficiency balance; summary curves are in Figure 4 (Right). (i) Sliding-window length $H$. Enlarging $H$ from 2 to 4 yields a clear accuracy gain (e.g., 94.36% $\to$ 95.10%), while H=5H{=}5 brings only marginal benefit with higher overhead (accuracy $\approx$95.05%, latency 101% $\to$ 110%, extra cost 3.5% $\to$ 6.7%). (ii) Parsimony weight $\lambda_{\mathrm{sw}}$. With H=3H{=}3, increasing $\lambda_{\mathrm{sw}}$ from 0.2 to 0.8 makes the policy thriftier (extra cost 5.6% $\to$ 3.3%; latency 106% $\to$ 100%) at a modest accuracy trade-off (94.85% $\to$ 94.38%). We default to $\lambda_{\mathrm{sw}}{=}0.4$ and H=3H{=}3.
@@ 6 Conclusion
In this paper, we present AgentAsk, a lightweight, architecture-agnostic clarifier that operates at the edge level of multi-agent chains to arrest error propagation with minimal, targeted questions. Grounded in a four-type taxonomy, AgentAsk couples supervised distillation with E-GRPO that aligns local fixes with end-task success under latency and cost constraints. Across multiple tasks, AgentAsk consistently improves performance over public multi-agent system methods while approaching large models’ performance at a fraction of the overhead. Beyond empirical gains, our edge-level perspective offers a principled recipe for reliable orchestration that complements role design, search, and self-checking. Future work spans uncertainty-aware ask gating, theory for edge-level interventions, tool- and human-in-the-loop integration, and robustness under adversarial or shifted settings.
@@
AgentAsk currently inherits much of its judgment quality from the underlying LLMs used as the teacher (for distillation) and as the lightweight clarifier backbone. As a result, its reliability and granularity of clarifications scale with these base models; under weaker backbones, gains may diminish even though the logic of edge-level intervention remains sound. Future work will explore model-agnostic uncertainty signals and hybrid features to reduce dependence on specific LLM strengths.
Appendix A Algorithm Workflow
This appendix concisely instantiates the edge controller in Eq. (2) under the constraint in Eq. (3). The details can be seen in Section 4.
Algorithm 1 AgentAsk: SFT & E-GRPO Training
@@ Appendix A Algorithm Workflow
2:Stage A: Build SFT Corpus
@@
9:Stage B: Supervised Fine-tuning (SFT)
@@
13:  Evaluate $\mathcal{L}_{\text{type}}$ and $\mathcal{L}_{\text{ask}}$; set $\mathcal{L}_{\text{SFT}}=\mathcal{L}_{\text{type}}+\lambda_{\text{ask}}\mathcal{L}_{\text{ask}}$
14:  $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}_{\text{SFT}}$
@@
16:Stage C: RL with E-GRPO
@@
25:     Sample $G$ candidates at $x_{t}$; score by $r_{t}^{\mathrm{edge}}$
26:     Form local advantages $A_{t}^{\mathrm{loc}}$; update with ratios $\rho_{t}$ (no $R$)

### v2

@@ 4.3 Clarification Optimization with Reinforcement Learning
Previous training enables us to transfer edge-level judgment and question quality from a strong evaluator. However, imitation learning has its limits: it only captures behavior from logged cases and cannot learn when not to ask, nor can it balance accuracy against constraints in new scenarios. The ask gate and addressee are discrete decisions whose true value emerges only after downstream effects, so credit must propagate through the entire sequence. Therefore, we continue with reinforcement learning to adapt the $\pi_{\theta}$ to the constrained objective in Equation (1), using shaped rewards that convert these trade-offs into learnable signals.
@@
Here, $n_{t+1}$ represents whether need to clarify again, that AgentAsk may ask multiple times at one edge.
Second, we enforce parsimony to keep clarifications scarce via a sliding counter $c_{t}$:
@@
To stabilize outputs and preserve brevity, we introduce a format bonus:
@@
These signals are aggregated into an edge reward that embodies our inspirations from taxonomy,
@@
where $s$ is the final success score of the task.
Optimization: E-GRPO.
We extend the principles of Group Relative Policy Optimization (GRPO) (DeepSeek-AI et al., 2025) by introducing a gradual global credit mechanism to be our E-GRPO. While GRPO optimizes policies based on edge-level rewards, we ensure that E-GRPO aligns local decisions with overall success by incorporating global feedback in a controlled manner.
For each edge $t$, we form a local advantage $A_{t}^{\mathrm{loc}}=r_{t}^{\mathrm{edge}}$ by the controller. The system-level signal $A_{t}^{\mathrm{glob}}=w_{t}\,(R-b)$ is distributed at time $t$ with nonnegative weights $w_{t}$ and an action-independent baseline $b$. The weight $w_{t}$ increases as the training progresses, ensuring that global feedback starts to influence the policy more as the agent gains confidence. Initially, the agent learns from edge-level rewards $r_{t}^{\mathrm{edge}}$, which guide the local decisions. However, to prevent the agent from getting trapped in local optima, we introduce global credit to reflect the overall success of the task. This global credit is gradually integrated into the learning process through a time-varying weight $\lambda_{R}(t)$, which increases as the training progresses.
@@
When $R$ is not available, the second term in Eq. (8) is simply absent, and training proceeds purely on edge-level selection; when $R$ arrives, it is injected uniformly through $A_{t}^{\mathrm{glob}}$, ensuring that minimal, successful clarifications remain aligned with end-to-end correctness while updates stay conservative through clipping and reference KL regularization.
This approach ensures that AgentAsk avoids becoming trapped in local optima, while progressively aligning edge-level decisions with global task outcomes, thereby improving the reliability of multi-agent systems by preventing early error propagation with minimal overhead.
@@
Table 1: Accuracy/Pass@1 across frameworks under three settings (origin, +GPT-5, +AgentAsk) on five benchmarks. For improved settings (+GPT-5 and +AgentAsk), bottom-right overlays show standard deviation in gray. Average column shows overall percentage-point gains (green $\uparrow$) relative to each framework’s origin.
@@ 5 Experiments
We assess whether a lightweight, edge-level clarifier improves reliability under unchanged orchestration, how it trades accuracy against latency and cost, and why these effects arise. Our study is structured around four research questions (RQ):
  • RQ1: Paradigm validity and end-to-end effectiveness. Under unchanged orchestration, does inserting an edge-level clarifier consistently improve end-task performance across frameworks and benchmarks?
  • RQ2: Efficiency under practical budgets. Can AgentAsk achieve these gains with low overhead in latency and extra cost, and how does it compare to the +GPT-5 upper-bound clarifier in the accuracy–efficiency trade-off?
  • RQ3: Ablations and robustness. Which components are necessary for the observed gains, and how stable is performance under key design choices and control parameters?
  • RQ4: Taxonomy-grounded mechanisms. Which error types drive most interventions and gains, and how does the edge-level clarification improve efficiency while minimizing unnecessary overhead?
@@ 5.1 Experimental Setup
Benchmarks & Baselines.
We evaluate AgentAsk on five public benchmarks spanning math reasoning (GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021b)), general-knowledge QA (MMLU (Hendrycks et al., 2021a)), and code generation (HumanEval (Chen et al., 2024a), MBPP (Austin et al., 2021)). Baselines include single-model prompting (IO, CoT, Self-Refine (Wei et al., 2022; Madaan et al., 2023)) and four representative multi-agent frameworks (GPTSwarm (Zhuge et al., 2024), AFlow (Zhang et al., 2025b), MaAS Zhang et al. (2025a), MasRouter (Yue et al., 2025)).
Implementation Details.
For each framework, we consider three settings that isolate the effect of the clarifier while holding orchestration constant: (i) origin uses the public implementation; (ii) +GPT-5 substitutes the clarifier with GPT-5 (OpenAI, 2025); (iii) +AgentAsk plugs in our lightweight module. The LLM executor is GPT-4o-mini-0718 (OpenAI, 2024); +AgentAsk uses Qwen-3-4B (Yang et al., 2025) as the backbone. Accuracy for GSM8K, MATH, and MMLU is reported; Pass@1 for HumanEval and MBPP. Due to baseline performance differences across frameworks, we normalize latency and extra cost by setting the origin framework to 100, and compare the relative increases for each configuration. Prompts are provided in the Appendix F. To better demonstrate our plug-and-play capabilities, we trained AgentAsk on one framework and then gradually integrated it into other frameworks.
@@ 5.2 RQ1: Effectiveness under Fixed Orchestration
Observation. Edge-level clarification consistently improves end-task performance under unchanged orchestration, suggesting that many failures originate at inter-agent handoff boundaries and can be mitigated by clarifying exchanged messages.
Results and analysis. We compare our lightweight AgentAsk with a strong clarifier, +GPT-5, which serves as an upper bound for the edge-level clarification paradigm. Table 1 shows that adding an edge-level clarifier yields consistent gains across frameworks and benchmarks. For each framework, both +AgentAsk and +GPT-5 improve performance over origin on all five datasets, and the average score increases by a clear margin. In particular, +AgentAsk improves the framework-level averages by about +3.29 to +3.45 points, while +GPT-5 improves them by about +4.40 to +4.69 points. The gains are consistent across diverse task types, including math, general knowledge, and code generation, indicating that edge-level issues are a common source of failure that can be mitigated by clarifying exchanged messages rather than altering the orchestration. As expected, +GPT-5 produces the largest improvements, supporting its role as an approximate upper bound for the clarification paradigm under the same pipeline. Notably, +AgentAsk recovers a substantial portion of this upper-bound gain with a lightweight learned module. The remaining gap is primarily attributable to efficiency considerations, which we analyze in RQ2.
To verify that the improvements are not specific to one decoding configuration, we additionally evaluate the main settings under multiple decoding temperatures and evaluation seeds, with full robustness results reported in Appendix C.1.
  GSM8K
Settings Accuracy Latency Extra Cost
@@
Table 2: Results of experiments on different types or training stages of clarifier with MasRouter@GSM8K.
Figure 4: (Left) Pareto frontier (x=latency, y=extra cost; bubble size=Acc), showing +AgentAsk nearing +GPT-5 at much lower overhead. (Middle) Error-Type distributions (DG/SC/RD/CG) across datasets. (Right) Resolved@Edge of four error types across different datasets in AgentAsk.
5.3 RQ2: Efficiency under Practical Budgets
Observation. AgentAsk attains most of the improvement with lower latency and extra cost, making it more efficient than heavy clarifiers.
Results and analysis. We quantify the accuracy–efficiency trade-off on MasRouter@GSM8K in Table 2. Using +GPT-5 yields the highest accuracy at 95.10, but increases latency to 129 and extra cost to 34.0. In contrast, AgentAsk with Qwen-3-4B achieves 94.86 accuracy with only 108 latency and 4.9 extra cost, reducing latency by 21 points and extra cost by 29.1 points while leaving a 0.24 point accuracy gap. The Llama-3.2-3B variant shows the same trend, further indicating that the efficiency advantage is not tied to a specific backbone.
Figure 4 (Left) shows that +AgentAsk stays near the accuracy envelope with low overhead, while +GPT-5 achieves slightly higher accuracy at much higher latency and cost. These results confirm that our edge-level clarifier is efficient under tight budgets, with the gap to the upper bound primarily due to capacity limitations rather than orchestration changes. We attribute this efficiency to the training objective that penalizes unnecessary interventions, encouraging the clarifier to ask only when the expected downstream benefit outweighs the overhead. Additional breakdowns across frameworks and datasets are reported in Appendix C.2.
@@ 5.4 RQ3: Mechanism and Robustness at Edge
Observation. The effectiveness of AgentAsk is driven by its ability to intervene at communication handoffs, improving accuracy while maintaining efficiency. Its robustness is confirmed through ablation studies and sensitivity analyses.
Results and analysis. Table 3 presents the results of ablation experiments on the reward components for AgentAsk with the MBPP benchmark. Removing individual reward components leads to a decrease in accuracy and an increase in latency and cost, showing that each component contributes to balancing efficiency and performance. The removal of the $r^{par}$ increases the latency and extra cost without much improvement in accuracy, emphasizing the importance of minimizing unnecessary clarifications for maintaining efficiency. Finally, the full set of results, including further details on ablation studies and experiments with PPO, GRPO, and our method, can be found in Appendix C.3, where Figure 5 demonstrates the sensitivity.
  MBPP
Settings Accuracy Latency Extra Cost
AgentAsk 83.55 103 7.7
[…]
Table 3: Results of ablation experiments on different reward settings with MasRouter@MBPP(red $\downarrow$ represents the decrease to each framework’s AgentAsk).
5.5 RQ4: Taxonomy Mechanisms at Edge
Observation. The performance of AgentAsk is influenced by the types of errors it addresses, with varying effectiveness across different error types.
Results and analysis. Figure 4 (Middle and Right) shows the distribution of error types across datasets, along with the Resolved@Edge percentages for each error type. Resolved@Edge represents the percentage of clarifications that successfully resolve the error without further intervention (illustrated in Eq 3). AgentAsk performs particularly well in tasks dominated by Data Gap and Signal Corruption errors. These errors are typically resolved with a single clarification, leading to high Resolved@Edge rates, 82.5% for DG and 75.4% for SC. DG makes up 36.8% and SC 30.9% of the errors across datasets, representing the majority of the issues encountered. For tasks involving more Referential Drift and Capability Gap, such as MATH and MMLU, the improvements are more limited. These errors are less frequent, with RD accounting for 20.7% and CG 11.6% of the errors, but they are more difficult to resolve with a single clarification. As a result, their Resolved@Edge rates are lower, 58.3% for RD and 49.5% for CG.
Overall, AgentAsk is highly effective in resolving simpler edge-level errors. However, for tasks with more complex errors like RD and CG, the improvements are more modest, as these errors stems from the nature of the hallucinations. For further analysis and additional results including qualitative case studies, please refer to Appendix D.
@@ 6 Conclusion
In this work, we provide novel insights into error propagation in multi-agent systems, emphasizing the critical role of communication at agent handoffs. Our error taxonomy identifies four common error types that drive system failures, laying the groundwork for effective intervention. Building on this taxonomy, we propose AgentAsk, a lightweight, architecture-agnostic clarifier that minimizes intervention by asking targeted questions at key communication points. This approach enables error containment with minimal overhead, achieving near large-model performance. Future work will explore uncertainty‑aware gating, theoretical analysis of edge‑level interventions, and integration with tools and human feedback.
@@
Despite its effectiveness in mitigating many forms of error propagation, AgentAsk cannot fully eliminate all errors within multi-agent systems. Certain errors originate from internal flaws within individual agents, particularly those tied to well-known limitations in large models, such as hallucinations and inconsistencies. These issues lie beyond the scope of a clarification module like AgentAsk and can only be addressed through more fundamental advancements in the capabilities of large models. Future research will need to focus on overcoming these intrinsic model limitations, with the hope that improvements in model robustness and accuracy will enable more reliable error resolution across the system.
Appendix A Supplementary Design Explaination
A.1 Modeling Details
We elaborate on the controller framework referenced in Section 4.1.
Edge State.
Each input produces a message graph $G=(V,E)$ unrolled into an edge sequence $\{e_{t}\}_{t=1}^{T}$, where each $e_{t}=(u_{t}\rightarrow v_{t})$ denotes a message handoff. The controller observes a local state at each step:
$x_{t}=\left(x^{\mathrm{in}}_{t},\;u_{t},\;v_{t},\;m_{t},\;h_{t}\right),$ (9)
where $x^{\mathrm{in}}_{t}$ is the user query or subgoal, $m_{t}$ is the message from $u_{t}$, and $h_{t}$ encodes recent local interaction context.
Controller Action.
Based on $x_{t}$, the controller selects an action:
$a_{t}=(z_{t},\tilde{v}_{t},q_{t}),$ (10)
where $z_{t}$ is a binary gate indicating whether to ask, $\tilde{v}_{t}$ selects which agent to query, and $q_{t}$ is the generated clarification question.
Policy Structure.
We define the controller policy as:
$\pi_{\theta}(a_{t}\mid x_{t})=\pi_{\theta}(z_{t}\mid x_{t})\;\left[\pi_{\theta}(\tilde{v}_{t}\mid x_{t})\;\pi_{\theta}(q_{t}\mid x_{t})\right]^{z_{t}}.$ (11)
Trajectory Modeling.
The full input interaction produces a trajectory:
$\tau=\left\{(x_{t},a_{t},r_{t})\right\}_{t=1}^{T},$ (12)
where $r_{t}$ is a response returned when $z_{t}=1$. The objective from Section 4.1 is instantiated here with:
  • $U(\tau)$ rewarding task success and helpful clarifications,
  • $C(\tau)$ measuring token and latency costs,
  • $B$ as the intervention budget.
This setup supports both imitation-based initialization and constrained reinforcement learning (see Appendix A.2).
Algorithm 1 AgentAsk Training
@@ A.1 Modeling Details
2:Stage A: Build Corpus
@@
9:Stage B: Learning from Demonstrations
@@
13:  Evaluate $\mathcal{L}_{\text{type}}$ and $\mathcal{L}_{\text{ask}}$; set $\mathcal{L}_{\text{Total}}=\mathcal{L}_{\text{type}}+\lambda_{\text{ask}}\mathcal{L}_{\text{ask}}$
14:  $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}_{\text{Total}}$
@@
16:Stage C:Clarification Optimization with RL
@@
25:     Sample $x_{t}$; score by $r_{t}^{\mathrm{edge}}$
26:     Form local advantages $A_{t}^{\mathrm{loc}}$; update with ratios $\rho_{t}$